An axle with a diameter of 100mm has a force of 10N applied to it, what is the torque at the wheel?


1
Expert's answer
2021-08-24T14:42:24-0400

"d = 100mm = 0.1m"

"r = \frac{d}{2}= 0.05m"

"F = 10N"

"\tau = Fr = 10*0.05 = 0.5N*m"


"\text{Answer: }\tau = 0.5N*m"